### v1

# 3DSRBench: A Comprehensive 3D Spatial Reasoning Benchmark

Abstract
3D spatial reasoning is the ability to analyze and interpret the positions, orientations, and spatial relationships of objects within the 3D space. This allows models to develop a comprehensive understanding of the 3D scene, enabling their applicability to a broader range of areas, such as autonomous navigation, robotics, and AR/VR. While large multi-modal models (LMMs) have achieved remarkable progress in a wide range of image and video understanding tasks, their capabilities to perform 3D spatial reasoning on diverse natural images are less studied. In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, with 2,772 manually annotated visual question-answer pairs across 12 question types. We conduct robust and thorough evaluation of 3D spatial reasoning capabilities by balancing the data distribution and adopting a novel FlipEval strategy. To further study the robustness of 3D spatial reasoning w.r.t. camera 3D viewpoints, our 3DSRBench includes two subsets with 3D spatial reasoning questions on paired images with common and uncommon viewpoints. We benchmark a wide range of open-sourced and proprietary LMMs, uncovering their limitations in various aspects of 3D awareness, such as height, orientation, location, and multi-object reasoning, as well as their degraded performance on images with uncommon camera viewpoints. Our 3DSRBench provide valuable findings and insights about the future development of LMMs with strong 3D reasoning capabilities. Our project page and dataset is available here.
1 Introduction
Recent large multi-modal models (LMMs) [1, 43, 3] have achieved significant improvements in a wide range of image and video understanding tasks, such as image captioning [2, 28], visual question answering [24, 20, 31, 10, 44], visual grounding [49], decision-making [34, 26, 7], and action recognition [48, 35]. However, recent studies have shown that even state-of-the-art LMMs exhibited limited 3D awareness [17, 36] and understanding of spatial relationships [46, 47], which are crucial for LMMs to develop a comprehensive grasp of the 3D scene, to enable high-level reasoning and interaction, and eventually, to achieve human-level intelligence.
Figure 1: 3D spatial reasoning performance comparison across leading models and human evaluation on 3DSRBench. Leading models (GPT-4o, Gemini-Pro, LLaVA-NeXT and Cambrian) lag behind human performance by at least -56% on average across four types of spatial relations (height, location, orientation, and multi-object). All models struggle with questions involving 3D orientation (e.g., “on the left”) and complex spatial reasoning (e.g., “viewpoint towards object”). See Fig. 2(a) for examples.
Figure 2: Overview of the our 3DSRBench. (a) Example questions from the four main types of 3D spatial reasoning questions, i.e., height, location, orientation, and multi-object reasoning. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy to remove left/right biases in 3D with paired VQAs (see Sec. 3.4).
To study the 3D spatial reasoning capabilities of LMMs, previous works exploited synthetic environments and generated images with 3D ground-truths [46, 47]. Visual question-answer pairs were automatically synthesized by applying pre-defined rules on the known 3D scene graphs and other object attributes. The synthetic images exhibit a significant domain gap with natural images and lacked the diversity and richness of contents in real-world. More recent works [13] explored real datasets with 3D annotations, e.g., Omni3D [6]. However, images in these datasets are still limited to specific domains, such as indoor rooms and self-driving scenes. In general, visual question-answer pairs generated with rule-based methods from 3D annotations (i) limit the scope of theirs datasets to a very small set of rigid objects, and (ii) cannot enable a fine-grained and robust evaluation of 3D spatial relationships that can only be achieved with human annotated datasets (see Sec. 3.1).
In this work we present the first comprehensive 3D spatial reasoning benchmark, 3DSRBench, that features a wide variety of 3D spatial reasoning questions on diverse and open-vocabulary entities, including rigid objects, humans, animals, and implicit concepts, such as logo on a car or arrow on a billboard. We manually annotate 2,100 visual question-answer pairs on natural images from the MS-COCO dataset [29], covering 12 subtypes of questions from 4 main categories, i.e., height, location, orientation, and multi-object reasoning. Each category of questions focus on different combinations of 3D properties, such as object 3D location, 3D ground plane, camera extrinsic calibration, and/or object 3D poses. Examples from each question category are presented in Figure 2(a).
Another challenge of 3D spatial reasoning arises from the 6D viewpoint of the camera, i.e., the 3D location and 3D orientation from which we are viewing the 3D scene. As shown in Fig. 4, 3D spatial reasoning questions can be easier for “common” 6D viewpoints, e.g., ones positioned at the eye level with natural viewing angles, while being more challenging for other “uncommon” viewpoints. Although uncommon viewpoints are less populated in most image datasets, cameras in embodied AI and robotics are often positioned in these uncommon viewpoints. Hence it is of crucial importance for LMMs to retain good 3D spatial reasoning performance for both common and uncommon viewpoints. To fairly compare the 3D spatial reasoning capabilities of LMMs w.r.t. different camera viewpoints, we annotate another 672 visual question-answer pairs on multi-view synthetic images rendered from the HSSD dataset [25].
Besides benchmarking a wide variety of open-sourced and proprietary LMMs, our 3DSRBench serves as an important diagnosis benchmark for developing 3D spatially intelligent LMMs. Inspired by previous studies on 3D awareness of visual foundation models [17, 36], our 3DSRBench takes one step further and evaluates LMMs on fundamental 3D spatial reasoning questions, which provide valuable insights regarding the 3D awareness of visual encoders [39, 11, 21, 38, 27] and the 3D reasoning capabilities of language models [14, 50, 45, 16]. Such results would shed light on downstream tasks that build on 3D spatial reasoning, such as automatic navigation and robotic manipulation.
To enable a comprehensive and robust evaluation of 3D spatial reasoning capabilities, our 3DSRBench adopts several key designs: (1) a balanced data distribution in multiple aspects, such as a roughly same number of yes/no answers and complementary images pairs that lead to opposite answers given the same question (see Fig. 2(b)); (2) avoiding questions with trivial answers; and (3) a novel FlipEval strategy designed for robust evaluation of 3D spatial reasoning questions.
In summary, our 3DSRBench significantly advances the evaluation of 3D spatial reasoning capabilities by manually annotating 2,100 visual question-answer pairs on natural images from MS-COCO [29] and 672 on multi-view synthetic images. We annotate diverse and open-vocabulary entities and adopt key designs to enable a robust and thorough evaluation that cannot be done with existing datasets. We benchmark a wide variety of open-sourced and proprietary LMMs and study their 3D spatial reasoning capabilities w.r.t. different types of 3D awareness. We further analyze the robustness of 3D spatial reasoning capabilities w.r.t. common and uncommon camera 6D viewpoints and study a range of LMMs with different visual encoder designs. Experimental results on different splits of our 3DSRBench provide valuable findings and insights that will benefit future research on 3D spatially intelligent LMMs.
2 Related Works
3D awareness of visual foundation models.
With the recent advancements in large multi-modal models [31, 30, 32], there has been a rising interest in applying these LMMs to a broader range of tasks, such as chatting about human poses [18], embodied question answering [37], and robotic manipulation [23, 22]. Notably, these tasks involve reasoning and interacting with the 3D scenes, which largely builds on the 3D awareness of vision encoders. Previous works studied the 3D awareness of visual foundation models by adopting proxy tasks, such as part correspondence [17] and pose estimation [36], and quantitatively evaluating the 3D awareness with linear probing. Our work can be considered as one step further — studying the 3D recognition and reasoning capabilities of LMMs by benchmarking their performance on fundamental 3D spatial relationship questions. Future research on downstream tasks, such as automatic navigation and robotic manipulation, could refer to the findings in our 3DSRBench and adopt LMMs with better 3D spatial reasoning capabilities.
3D spatial reasoning.
Previous studies on 3D spatial reasoning capabilities of vision-language models often adopt synthetic environments, e.g., Blender, with controllable simulation and 3D groundtruths for automatic question-answer generation [12, 46, 47]. However, synthetic images in these datasets exhibit a large domain gap with natural images and it remains unclear if insights and findings from these datasets would generalize to the real image domain. More recent works, such as SpatialRGPT [13] and Cambrian-1 [44], built on existing datasets with 3D annotations [6, 42, 5, 19, 41, 8] and generated visual question-answer pairs with pre-defined rules. Despite the improved image qualities, they are essentially limited to a small number of rigid object categories in Omni3D [6]. To enable a comprehensive and valuable evaluation of the 3D spatial reasoning capabilities, we manually annotate visual question-answer pairs, with diverse and open-vocabulary entities, such as logos on a car or arrows on the billboard, and questions about a comprehensive list of 3D spatial relationships from four main categories, i.e., height, location, orientation, and multi-object reasoning.
3 3DSRBench
In this section we introduce the 3DSRBench for comprehensively analyzing the 3D spatial reasoning capabilities of large multi-modal models (LMMs). We start by introducing the design considerations of 3DSRBench in Sec. 3.1, specifically how these design choices lead to a robust and valuable evaluation of 3D spatial reasoning capabilities. Then we present the four main question types in Sec. 3.2, as well as the challenges in each type of questions. Next we discuss the three splits of our 3DSRBench and their scopes in Sec. 3.3. In Sec. 3.4 we present our evaluation strategies, including CircularEval and FlipEval. Lastly in Sec. 3.5 we provide an overview of our data collection procedures and the summary statistics.
3.1 Design of 3DSRBench
When developing 3DSRBench, we incorporate the following four key designs to enable a robust and valuable evaluation of 3D spatial reasoning capabilities. First, our 3D spatial reasoning questions are based on open-vocabulary entities. Previous spatial reasoning benchmarks [9, 44] largely relied on existing datasets with 3D annotations [6], which limited their scope to a small number of rigid object categories. In our 3DSRBench, we annotate 3D spatial reasoning questions across a broad range of open-vocabulary entities (see Fig. 2), enabling a thorough analysis of the 3D awareness and 3D reasoning capabilities of LMMs over diverse, commonly encountered real-world objects. Next, we avoid questions with trivial answers. For instance, when comparing which of the two objects has a smaller 3D distance to a third anchor object, we avoid the cases when there is a significant gap between the two distances. We focus on samples where the two distances are close, but the correct answer can be derived with annotators’ consensus (see Sec. 3.5). Moreover, we implement a balanced data distribution in various aspects, such as a roughly same number of yes/no answers and complementary image pairs [20] that lead to opposite answers given the same 3D spatial reasoning question (see Fig. 2(b)). This effectively removes priors in the answer distribution, e.g., pedestrians are often located lower than street lights, or the fact that objects higher in 3D space are also higher in 2D image plane. This design ensures that models cannot exploit biases or shortcuts for a higher benchmark performance. Lastly, we adopt special evaluation strategies for robust evaluation, including previous CircularEval [33] and our novel FlipEval (see Sec. 3.4).
3.2 Question Types
(a) Height questions with different camera pitch rotations.
(b) Comparison between 2D and 3D spatial reasoning questions.
Figure 3: Challenges of 3D spatial reasoning questions in our 3DSRBench. See Sec. 3.2. (a) Height questions requires 3D spatial reasoning over a combination of camera extrinsics and object 3D locations. Notice how different camera pitch rotations play a crucial role to determine the final answer. (b) Previous 2D spatial reasoning questions can be addressed by analyzing objects’ 2D locations and depths, while our orientation questions require complex 3D spatial reasoning on objects’ 3D orientations and 3D locations.
We present the 4 types of 3D spatial reasoning questions in our 3DSRBench. We discuss why they are challenging for LMMs and what kinds of 3D awareness and 3D spatial reasoning are needed to succeed in each type of questions. We present an overview of the 4 question types in Tab. 1.
Height questions.
For height-related question, we study if models can determine which of the two given objects is positioned higher in the 3D world space. To correctly answer the questions, a model must (i) calibrate camera extrinsics, such as roll and pitch rotations, and then (ii) detect 3D locations of the objects in the 3D world space. This task poses a significant challenge for large multi-modal models as these fine-grained 3D knowledge are hard to derive from the weak language supervision in standard multi-modal pre-training. In Figure 3(a) we illustrate two examples of height questions. Notice how different pitch rotations of the camera, i.e., viewing from above in the left figure and viewing upward in the right figure, play a crucial role to determine the final answer. In both examples, relying solely on the 2D locations within the image plane or the 3D locations in the camera coordinate system would lead to incorrect answers.
Location questions.
There are three subtypes of location-related questions, i.e., determining (i) if two objects are next to or far from each other, (ii) which of the two objects is closer to the camera, and (iii) if an object is directly above or underneath another object. Models must not only ground the 2D locations of the objects, but also understand the depth of field presented in the image. Consider the location question in Fig. 2(a). Although the 2D locations of the man and the hydrant are close, they are in fact far away from each other in the 3D space. Humans can determine the answer by estimating a rough depths of the two objects, or from other visual cues, such as how the pedestrian walk leads towards the vanishing point. Other examples include the top two questions in Fig. 2(b), which also require an understanding of the depth field.
Orientation questions.
Orientation-related questions study the 3D spatial reasoning that involves estimating the 3D orientation of an object. These questions are divided into three subtypes: determining which “side” of an object faces the camera, whether an object is in front of or behind another, and if an object is positioned on the left or right side of another. Unlike previous 2D spatial reasoning questions [9] that focus on spatial relationships w.r.t. the viewer’s perspective, our orientation-related questions emphasize spatial relationships from the object’s perspective. As demonstrated in Fig. 3(b), 2D spatial reasoning questions can be addressed by analyzing objects’ 2D locations and depths. Meanwhile, our orientation questions require estimating objects’ 3D orientation and perform 3D spatial reasoning across various dimensions of 3D information.
Multi-object reasoning questions.
Multi-object reasoning questions consider the 3D spatial relationships between multiple objects, such as asking which side of an object is facing another object, or with three objects, asking which of the given objects is facing towards or closer to the third object. In general, these questions require more advanced 3D awareness than simpler 3D concepts such as “closer” (to the camera) or “higher”, and require more complex reasoning over multiple objects’ 3D locations and orientations.
Type # Subtypes Camera Loc. Orient. Reasoning
Height 1 ++
Location 3 ++
Orientation 3 ++
Multi-Object 5 ++
Table 1: Overview of the 4 types of 3D spatial reasoning questions. Each type of 3D spatial reasoning questions require different combinations of 3D awareness, i.e., camera calibration, 3D location, 3D orientation, and multi-object reasoning.
3.3 Benchmark Splits
(a) Orientation questions on multi-view images from common (left) and uncommon (right) camera 6D viewpoints.
(b) Multi-object reasoning questions on multi-view images from common (left) and uncommon (right) camera 6D viewpoints.
Figure 4: Examples of the paired visual question-answer data in our 3DSRBench-synthetic. (a) Example questions from the four main types of 3D spatial reasoning questions. (b) To enable a robust evaluation of the 3D spatial reasoning capabilities, we collect complementary images that lead to opposite answers given the same question and adopt a novel FlipEval strategy (see Sec. 3.4).
Our 3DSRBench is composed of three splits, a real split with 2,100 3D spatial reasoning questions on MS-COCO images [29] and two synthetic splits with 672 questions on synthetic images rendered with 3D scenes in HSSD [25]. We evaluate the standard 3D spatial reasoning capabilities of LMMs on visual question-pairs from the real split, and with the synthetic split, we study the robustness of 3D spatial capabilities w.r.t. common and uncommon camera 6D viewpoints by analyzing the gap between the synthetic-common and synthetic-uncommon splits.
With the HSSD 3D scenes and controllable photorealistic rendering, we obtain multi-view images of the same 3D scene, each rendered with a common and an uncommon viewpoint. We ask the same 3D spatial reasoning question regarding the two images and study if models can obtain the correct answers on common and uncommon camera 6D viewpoints. We define “common” viewpoints as 6D camera poses with zero roll rotation, small pitch rotation, and taken from the height of a human, simulating the typical perspective when people take pictures. Conversely, “uncommon” viewpoints include 6D poses with noticeable roll rotation, large pitch rotation, or perspectives taken close to the ground or from a high location. The two synthetic splits are denoted by synthetic-common and synthetic-uncommon and examples from the two splits are demonstrated in Fig. 4. Notice how the answers by GPT-4o are correct when shown the image from a common camera 6D viewpoint and wrong when prompted from an uncommon viewpoint, despite both images present a clear view of the 3D scene and humans can derive the correct answers without any difficulty.
3.4 Evaluation
Since all 3D spatial reasoning questions in 3DSRBench have two or four answer choices, we formulate these questions as multiple choice questions with two or four options. To accommodate the free-form answers predicted by pretrained LMMs, we follow [33] and adopt LLM-involved choice extraction to obtain the predicted label. To enable a robust evaluation of various 3D spatial reasoning capabilities, we adopt the following two designs during testing:
CircularEval [33].
To avoid the bias of choice ordering and the influence of random guessing for multiple choice questions, we adopt CircularEval [33] for more robust benchmark performance. Specifically we feed each question into the LMM two or four times, each with a different ordering of the answer choices. The LMM is considered successful in answering this question only if the predicted answer is correct for all passes.
FlipEval.
We further propose FlipEval to remove left/right biases in 3D with paired visual question-answer pairs. By applying horizontal flip to the image, we obtain a new visual question. The answer would generally remain the same, such as for location and height questions, but when it involves 3D spatial relationships such as “left” and “right”, the answer would change. We illustrate this idea in Fig. 2(b), where the elephant logo is the on the left of the truck but changes to the right side after horizontal flipping. FlipEval effectively removes left/right biases in 3D spatial relationship, such as driver often sitting on the left side of the car or most people holding tools in their right hands. Lastly FlipEval also helps to avoid the influence of random guessing and enriches the image distribution in our 3DSRBench.
3.5 Data Collection
We employ three annotators to annotate a total of 2,772 unique visual question-answer pairs across 12 question types. We follow the annotation principles as discussed in Sec. 3.1 and adopt an two-stage pipeline to ensure various criteria are met. Specifically after annotations are collected in the first stage, we review the quality of the collected data and reject samples with low quality or ones that lead to imbalanced data distribution. Further new annotations are collected if necessary. Furthermore, we collect human responses for all visual question-answer pairs and disregard samples that don’t reach consensus by human annotators.
4 Experiments
In this section, we start by introducing our experimental settings in Sec. 4.1. Then we benchmark the 3D spatial reasoning capabilities of various open-sourced and proprietary LMMs on our 3DSRBench in Sec. 4.2. Next we consider variants of LMMs with different visual encoders and study how the 3D awareness of visual encoders contribute to the final 3D spatial reasoning capabilities in Sec. 4.3. In Sec. 4.4 we evaluate various LMMs on our 3DSRBench-synthetic and analyze the robustness of LMMs w.r.t. uncommon camera viewpoints. Lastly we present some failure cases of GPT-4o in Sec. 4.5, revealing some interesting findings about the 3D spatial reasoning of LMMs.
4.1 Experimental Settings
Testing data augmentation.
We develop rule-based methods to augment the unique annotated visual question-answer pairs and obtain a larger number of testing data with a balanced and rich set of 3D spatial relationships. For instance, given a question asking which object is higher in the 3D world space, we generate a new question asking which object is lower in the 3D world space. We end up with 2,625 questions on MS-COCO images [29], i.e., 3DSRBench-real, and 846 questions on synthetic images, i.e., 3DSRBench-synthetic.
4.2 Results on 3D Spatial Reasoning Capabilities
Model 3DSRBench-real
Overall Height Loc. Orient. Multi.
Baselines
Random 20.9 25.0 25.0 16.8 20.1
Random++ 45.8 50.0 50.0 41.7 45.0
Human 95.7 92.9 96.4 97.7 94.9
Open-sourced
LLaVA-v1.5-7B [30] 36.8 38.5 46.4 27.7 31.8
Cambrian-1-8B [44] 44.1 25.6 57.0 36.5 43.1
LLaVA-NeXT-8B [32] 49.6 50.6 62.7 36.8 43.6
Proprietary
Claude-Sonnet [43] 46.9 49.6 60.0 32.8 41.2
Gemini-1.5-Flash [43] 39.2 39.8 59.9 13.2 33.6
Gemini-1.5-Pro [40] 49.1 50.8 62.9 37.5 41.3
GPT-4o-mini [1] 39.1 42.1 51.8 23.4 34.6
GPT-4o [1] 45.3 49.4 62.3 23.0 40.1
Table 2: Experimental comparison of state-of-the-art large multi-modal models on our 3DSRBench. Results show that state-of-the-art LMMs exhibit limited 3D spatial reasoning capabilities. Please refer to Sec. 4.2 for detailed analyses.
3DSRBench
LLM Vision Encoder Connector Mean Height Loc. Orient. Multi.
Baseline
Vicuna-v1.5-7B [50, 30] CLIP-L14-336 [39] 2xMLP 36.8 38.5 46.4 27.7 31.8
Mixed Encoders
Vicuna-v1.5-7B [50, 30] CLIP-L14-336 [39] + DINOv2-L14-224 [38] 2xMLP 37.2 45.9 42.2 28.7 33.6
Vicuna-v1.5-7B [50, 30] CLIP-L14-336 [39] + MAE-H14 [21] 2xMLP 33.1 42.7 39.2 26.1 27.5
Vicuna-v1.5-7B [50, 30] CLIP-L14-336 [39] + SAM-L [27] 2xMLP 27.9 44.6 34.4 16.5 21.5
Connectors
Vicuna-v1.5-7B [50, 30] CLIP-L14-336 [39] + DINOv2-L14-224 [38] SVA [44] 37.8 46.0 43.1 26.5 35.9
Vicuna-v1.5-7B [50, 30] CLIP-L14-336 [39] + MAE-H14 [21] SVA [44] 34.1 45.3 38.6 25.3 30.2
Table 3: Experimental results on LMMs with various vision encoder setups. We use LLaVA-v1.5-7B as the baseline model and studies how vision encoders with different features contribute to the final 3D spatial reasoning abilities of LMMs.
Model 3DSRBench-synthetic-common 3DSRBench-synthetic-uncommon
Overall Height Loc. Orient. Multi. Overall Height Loc. Orient. Multi.
Random 20.9 25.0 25.0 16.8 20.1 20.9 25.0 25.0 16.8 20.1
Random++ 45.8 50.0 50.0 41.7 45.0 45.8 50.0 50.0 41.7 45.0
Open-sourced
LLaVA-v1.5-7B [31] 42.0 40.0 50.6 20.8 47.6 38.0 41.0 43.6 17.9 45.2
Cambrian-1-8B [44] 48.1 37.5 56.1 39.6 47.6 39.9 35.0 45.7 29.2 41.9
LLaVA-NeXT-8B [32] 45.5 65.0 57.9 10.4 50.0 36.8 47.5 44.5 7.3 46.0
Proprietary
Qwen-VL-Plus [4] 30.7 35.0 37.8 30.2 20.2 21.0 15.0 25.0 22.9 16.1
Qwen-VL-Max [4] 55.2 62.5 69.5 31.2 52.4 48.6 52.5 59.8 24.0 51.6
Claude-Sonnet [3] 47.4 47.5 58.5 26.0 49.2 39.4 60.0 48.2 16.7 38.7
Gemini-1.5-Flash [40] 44.6 57.5 59.8 13.5 44.4 37.7 42.5 45.7 11.5 46.0
Gemini-1.5-Pro [40] 59.9 65.0 69.5 50.0 53.2 49.5 42.5 52.4 40.6 54.8
GPT-4o-mini [1] 46.5 47.5 53.7 36.5 44.4 40.3 42.5 43.9 33.3 40.3
GPT-4o [1] 51.2 70.0 70.1 17.7 46.0 44.3 60.0 58.5 15.6 42.7
Table 4: Experimental results on our 3DSRBench-synthetic-common and 3DSRBench-synthetic-uncommon. We study the robustness of 3D spatial reasoning capabilities of LMMs by analyzing the performance gap between the two splits with images from the same 3D scene but from “common” and “uncommon” viewpoints. We find that LMMs does not generalize well to images with 6D camera viewpoints less represented in their training set. See Sec. 4.4 for detailed discussions.
We benchmark a wide range of open-sourced and proprietary LMMs on our 3DSRBench real and analyze their 3D spatial reasoning capabilities on different categories of questions. We consider two baseline results: (i) random: a simple baseline that predicts random answers for all visual questions. (ii) random++: a stronger random baseline that predicts consistent answers given different choice orders of a same visual question in CircularEval. We report the full results in Tab. 2. (iii) human: a human-level performance established by human evaluators that did not participate in the data annotation process. We report the full results in Tab. 2.
We make the following observations: (i) State-of-the-art LMMs have limited 3D spatial reasoning capabilities, as found by low performance achieved by state-of-the-art open-sourced and proprietary LMMs, falling far behind human-level performance. (ii) Scaling laws for LMMs are not effective for 3D spatial reasoning. Results show that despite significant more training data and computation spent on the proprietary LMMs, they do not show clear advantages over open-sourced counterparts, featuring high-quality data with efficient training setups. Standard scaling laws demonstrate diminishing returns for 3D spatial reasoning capabilities and we believe more effective approaches, e.g., 3D-aware data, architecture, and training designs, would be necessary to significantly advance 3D spatial reasoning.
4.3 3D Awareness of Vision Encoders
We further study the 3D awareness of vision encoders by benchmarking variants of LLaVA-v1.5-7B [31] on our 3DSRBench with various vision encoders. Specifically, we consider a range of pretrained visual foundation models, i.e., CLIP [39], MAE [21], DINOv2 [38], and SAM [27], and enhance the baseline model with mixed encoders. This can retain the performance on standard benchmarks with the semantic features and potentially involve useful features for 3D reasoning. Next we consider the spatial vision aggregator (SVA) [44] to further fuse the features from multiple vision encoders.
We quantitatively evaluate these LMM variants on our 3DSRBench and report the results in Tab. 3. Results show that with mixed encoders, DINOv2 can improve the overall 3D spatial reasoning abilities of LMM. However, we notice significant improvements for height questions when adopting MAE and SAM as vision encoder, suggesting that having richer visual features could help localize objects better. With SVA [44], we can further improve the LMM with mixed encoder from 37.2% to 37.8%, demonstrating that fusing the semantic features with 3D-aware features from DINOv2 would benefit subsequent reasoning.
4.4 Robustness to Uncommon Camera Viewpoints
Figure 5: Failure cases of GPT-4o on our 3DSRBench. We find that GPT-4o cannot perform rigorous 3D spatial reasoning and resort to visual cues for reasoning. The bottom two examples show that GPT-4o cannot perform certain complicated 3D spatial reasoning and would produce a short and succinct guess instead.
We study the robustness of 3D spatial reasoning capabilities w.r.t. common and uncommon viewpoints. We evaluate a variety of open-sourced and proprietary LMMs on our 3DSRBench-synthetic-common and 3DSRBench-synthetic-uncommon splits. As demonstrated by the results in Tab. 4, we can see that all LMMs exhibit significantly degraded performance when generalizing from “common” viewpoints to “uncommon” viewpoints, e.g., a 6.9% drop in accuracy for GPT-4o, a 10.4% drop for Gemini-Pro, and a 8.7% drop for LLaVA-NeXT-8B [32]. We visualize two failure examples of GPT-4o in Fig. 4, showing how GPT-4o is capable of predicting the correct answer when prompting with an image from a common viewpoint but fails when being asked about the same question but the image is rendered from an uncommon viewpoint of the exact same scene.
We attribute such degraded performance on “uncommon” viewpoints to two factors: (i) shift in camera 6D viewpoint distribution between training data and images in our 3DSRBench-synthetic-uncommon, and (ii) state-of-the-art LMMs adopt an implicit representation of 3D scenes. Current LMMs are heavily built on the scaling law of data-driven approaches. By leveraging large-scale data for pre-training, alignment, and instruction tuning, models develop an implicit 3D awareness and enable subsequent 3D spatial reasoning capabilities. Despite their success on a wide range of academic and empirical benchmarks, our results demonstrate that state-of-the-art LMMs, open-sourced or proprietary, face severe challenges generalizing to less represented data, which in our case, are images from uncommon camera 6D viewpoints. Notably, although images from these “uncommon” viewpoints are scarce in current datasets, they are in fact very common for specific downstream tasks, such as autonomous navigation and robotic manipulation, where the cameras are often placed on robots rather than held in hand by humans. Degraded performance of current LMMs would largely limit their potential in these areas.
4.5 Failure Cases
We present some failure cases of GPT-4o in Fig. 5. We make the following two observations: (1) GPT-4o cannot perform rigorous 3D spatial reasoning and resort to various visual cues for reasoning. For the top figure, GPT-4o fail to estimate a rough distance of the two objects, while relying on empirical facts that objects closer to the camera demonstrate more detailed appearances. For the second figure GPT-4o represent orientations with “facing toward approaching traffic”, which is a vague and ineffective orientation representation, leading to the false predicted answer. (2) GPT-4o cannot perform certain complicated 3D spatial reasoning. When being asked about complicated 3D spatial reasoning questions, e.g., ones that require 3D spatial reasoning over 3D locations and orientations, GPT-4o would give a short and succinct answer without complex reasoning. This can further serve as a criteria to study the decision making of GPT-4o, e.g., when GPT-4o skip reasoning and predict answer directly, and analyze its weakness.
5 Conclusions
In this work we study the 3D spatial reasoning capabilities of LMMs. We present a new benchmark, 3DSRBench, by manually annotating 2,100 visual question-answer pairs on MS-COCO images, featuring diverse and open-vocabulary entities, such as logos on a car or arrows on a billboard, and a balanced data distribution for robust evaluation. To study the robustness of 3D spatial reasoning capabilities w.r.t. camera 6D viewpoints, we further annotate 672 visual question-answer pairs on synthetic multi-view images, each with a common and an uncommon camera viewpoint. Our benchmark further serves as an important diagnosis benchmark for developing 3D spatially intelligent LMMs, shedding light on downstream tasks that require 3D spatial reasoning. Experimental results on different splits of our 3DSRBench provide valuable findings and insights that will benefit future research on 3D spatially intelligent LMMs.
Supplementary Material
A Baseline LMMs
Proprietary LMMs.
To analyze the 3D spatial reasoning capabilities of state-of-the-art LMMs, we explore a variety of proprietary LMMs, e.g., Qwen-VL, Claude, Gemini, and GPT. We follow Duan et al. [15] and formulate all questions as multiple choice questions with circular evaluation enabled.
LLaVA-v1.5 [30].
LLaVA-v1.5-7B is a strong open-sourced LMM baseline built on a Vicuna-v1.5 LLM and a CLIP-ViT-L vision encoder. It extends the visual instruction tuning framework [31] with an MLP connector and a scaled up image resolution.
Cambrian-1 [44].
Cambrian-1 is a strong LMM with vision-centric designs. It features an advanced connector design, spatial vision aggregator (SVA), and high-quality visual-instruction tuning data.
LMMs with various vision encoder designs.
We further experiment on a family of LMMs, extending the LLaVA-v1.5 baseline with various visual encoders design. We study the impact of the 3D awareness of visual encoders on the final 3D spatial reasoning capabilities. We adopt the same LLM and training strategy, exploring: (i) different mixed encoders: involving a second visual encoder besides CLIP, e.g., DINOv2 [38], MAE [21], and SAM [27]; and (ii) different visual connectors: a standard MLP connector and spatial vision aggregator (SVA) [44].
B Data Statistics
As discussed in Sec. 3, our 3DSRBench consists of three splits, a real split with 2,100 visual question-answer pairs manually annotated on diverse and open-vocabulary objects from the COCO dataset [29], and two synthetic splits synthetic-common and synthetic-uncommon with 336 visual question-answer pairs in each split, featuring same questions on a 3D scene rendered from a common and an uncommon viewpoint.
Our 3DSRBench consists of 12 question types from four main categories, as demonstrated in Figure 1. All question types are equally distributed, e.g., 175 questions for each question type in 3DSRBench-real.
To show that our 3DSRBench is balanced in answers, we visualize the distribution of closed-form answers in Fig. 6, i.e., one of the four directions (front, left, etc.) or yes/no answers rather than choosing from two given open-vocabulary entities.
C Example Questions in 3DSRBench
We present two example questions for each of the 12 question types in Fig. 7 (height and location questions), Fig. 8 (orientation questions), and Fig. 9 (multi-object reasoning questions).
Figure 6: Distribution of closed-form answers in our 3DSRBench.
Figure 7: Two example questions for each of the 12 question types (part I): height and location questions.
Figure 8: Two example questions for each of the 12 question types (part II): orientation questions.
Figure 9: Two example questions for each of the 12 question types (part III): multi-object reasoning questions.